Solve this equation by using the quadratic form.
x(1/2) + (sqrt 3)x(1/4) - 18 = 0
or
x to the 1/2 power plus sqrt of 3 x to the 1/4 power - 18 = 0.
Any help will be greatly appreciated.

Answer by stanbon(75887)

x(1/2) + (sqrt 3)x(1/4) - 18 = 0
----
Let w = x^(1/4)
Then w^2 = x^(1/2)
-----
Substitute to get:
w^2 + sqrt(3)w - 18 = 0
----
Quadratic Formula:
---
w = [-sqrt(3) +- sqrt(3-4*-18)]/2
---
w = [-sqrt(3) +- sqrt(75)]/2
---
w = [-sqrt(3)+-5sqrt(3)]/2
---
w = [-3sqrt(3)] or w = 2sqrt(3)
-------
Now, solve for "x" using Let w = x^(1/4)
----
Solve: -3sqrt(3) = x^(1/4)
x = [-3sqrt(3)]^4 = 81*9 = 729
----
Solve: 2sqrt(3) = x^(1/4)
x = [2sqrt(3)]^4 = 16*9 = 144
------
Check these answers in the original equation:
========
Cheers,
Stan H.